# Defending Multimodal Backdoored Models by Repulsive Visual Prompt Tuning

### v2

@@ Abstract
Multimodal contrastive learning models (e.g., CLIP) can learn high-quality representations from large-scale image-text datasets, while they exhibit significant vulnerabilities to backdoor attacks, raising serious safety concerns. In this paper, we reveal that CLIP’s vulnerabilities primarily stem from its tendency to encode features beyond in-dataset predictive patterns, compromising its visual feature resistivity to input perturbations. This makes its encoded features highly susceptible to being reshaped by backdoor triggers. To address this challenge, we propose Repulsive Visual Prompt Tuning (RVPT), a novel defense approach that employs deep visual prompt tuning with a specially designed feature-repelling loss. Specifically, RVPT adversarially repels the encoded features from deeper layers while optimizing the standard cross-entropy loss, ensuring that only predictive features in downstream tasks are encoded, thereby enhancing CLIP’s visual feature resistivity against input perturbations and mitigating its susceptibility to backdoor attacks. Unlike existing multimodal backdoor defense methods that typically require the availability of poisoned data or involve fine-tuning the entire model, RVPT leverages few-shot downstream clean samples and only tunes a small number of parameters. Empirical results demonstrate that RVPT tunes only 0.27% of the parameters in CLIP, yet it significantly outperforms state-of-the-art defense methods, reducing the attack success rate from 89.70% to 2.76% against the most advanced multimodal attacks on ImageNet and effectively generalizes its defensive capabilities across multiple datasets. Our code is available on https://github.com/zhangzf01/RVPT.
@@ 1 Introduction
Figure 1: Perturbation resistivity (PR) across different layers of the encoders under various perturbations, including the trigger pattern. A higher PR value indicates less sensitivity to input perturbation. Four encoder variants are evaluated: (a) ViT trained exclusively on clean ImageNet [44]. (b) ViT in backdoored CLIP [45]. (c) ViT in backdoored CLIP tuned with Visual Prompt Tuning (VPT) on ImageNet. (d) ViT of the backdoored CLIP tuned with our proposed RVPT on ImageNet. Specifically, The trigger of the backdoored CLIP is BadCLIP [31]. Detailed experimental settings and additional PR results for CLIP backdoored by other triggers can be found in the supplementary material.
@@
Recently, Contrastive Language-Image Pretraining (CLIP) [45] has emerged as a powerful base model across multiple domains. Unlike traditional models that rely on uni-modal supervision, CLIP aligns visual and textual signals to learn rich representations from numerous image-text pairs in the pre-training stage. This unique training paradigm enables its impressive performance in open-vocabulary visual recognition [57] and robustness to distribution shift [13].
@@
However, despite CLIP’s success in visual recognition, recent research [4, 5] has revealed its vulnerabilities against backdoor attacks in downstream tasks. The vulnerabilities can be roughly categorized into two aspects. First, concerns arise from its training data collection process: the data is noisy, uncurated, and sourced from the web, which facilitates adversarial poisoning of the image-text pairs. Second, CLIP demonstrates heightened susceptibility to a smaller number of poisoned training samples: recent studies [4] have empirically shown that poisoning CLIP may require injecting orders of magnitude fewer poisoned data into the training dataset than poisoning conventional supervised models. Once poisoned during pre-training, the backdoored CLIP will inevitably classify images with an adversarially designed trigger into the predefined target class.
@@
CLIP’s vulnerabilities to backdoor attacks have inspired the development of various defense methods. However, most of these methods necessitate fine-tuning the entire model’s parameters [3] or require the availability of poisoned data [61, 22, 60], which could be resource-intensive or even unrealistic. Thus, to achieve an efficient, effective, and practical defense, we are dedicated to developing a defense method based on visual prompt tuning (VPT) [24], utilizing only a small portion of downstream clean samples to encourage the model to ignore the trigger feature in the poisoned images. However, simply adopting VPT proves ineffective experimentally, as the trigger feature cannot be extracted from clean samples, preventing VPT from learning how to disregard them. Alternatively, VPT can help CLIP learn to extract in-dataset predictive features from clean samples using cross-entropy loss. If the visual encoder is guided to focus exclusively on these features, CLIP will gain backdoor robustness since the trigger features learned in pre-training are ignored during backdoor attacks.
@@
Nevertheless, we observe that CLIP’s encoded visual features, even after VPT, remain highly sensitive to perturbations, which suggests that it still encodes features beyond in-dataset predictive features, leaving it susceptible to potential backdoor attacks. To quantify this sensitivity, we introduce the Perturbation Resistivity (PR) metric, defined as the cosine similarity between the encoded feature $f^{v}(\boldsymbol{x})$ of the image $\boldsymbol{x}$ and the encoded feature $f^{v}(\boldsymbol{x}+\boldsymbol{\delta})$ of its perturbed version: $\boldsymbol{x}+\boldsymbol{\delta}$, where $f^{v}(\cdot)$ is the visual feature extractor:
$\text{PR}(\boldsymbol{x},\boldsymbol{\delta})=\cos(f^{v}(\boldsymbol{x}),f^{v}(\boldsymbol{x}+\boldsymbol{\delta})).$ (1)
A lower PR value indicates greater sensitivity of the encoded features to input perturbations, making the model more susceptible to backdoor triggers. To illustrate this vulnerability of CLIP, we randomly sample 1,000 images from ImageNet and apply various perturbations. We then compare the average PR values of these images across different model variants. As shown in Figure 1 (b) (c), the deeper visual layers of CLIP exhibit significantly higher sensitivity to input perturbations, even after VPT, which allows the trigger to maliciously reshape the encoded visual features layer by layer, enabling successful attacks. In contrast, Figure 1 (a) shows that the ViT supervised on clean ImageNet exhibits lower sensitivity, because it learns only from ImageNet, preventing it from encoding irrelevant features beyond its dataset. Therefore, to reduce CLIP’s sensitivity to input perturbations and enhance its backdoor robustness, we should suppress irrelevant features learned during pre-training and guide it to focus on encoding in-dataset predictive features.
@@
To address this challenge, we propose a novel approach termed Repulsive Visual Prompt Tuning (RVPT), which improves CLIP’s PR by allowing it to exclusively encode the predictive features in the downstream tasks in a few-shot manner. As illustrated in Figure 2, RVPT jointly optimizes two objectives: (1) a multi-layer feature-repelling loss that minimizes the cosine similarity between prompted features and original CLIP features at each layer, and (2) cross-entropy loss for downstream task accuracy. This mechanism forces the model to only extract features that contribute to CE loss optimization during training, while ignoring non-essential features (e.g., noise or trigger pattern). As a result, Figure 1 (d) shows that RVPT significantly improves the PR of CLIP on ImageNet, hence effectively defending against the state-of-the-art multimodal backdoor attack. Moreover, extensive experiments demonstrate our method’s effectiveness across multiple datasets and backdoor attacks.
@@ 2 Related Work
Backdoor attacks and defenses on supervised learning. Backdoor attacks have become an increasingly serious security issue because more and more practitioners choose to adopt third-party datasets, platforms, or even backbones to reduce costs. Research on backdoor attacks has primarily focused on designing triggers that enhance stealthiness [7, 50] and effectiveness [34, 40]. For backdoor defenses, researchers employ various strategies to mitigate backdoor attacks, including: (1) pre-processing defense [49] (2) pre-training defense [6] (3) post-training defense [68, 55] (4) test-time defense [15]. Specifically, our work aligns with the post-training defense category, which assumes that defenders do not have access to the poisoned training data and must eliminate the backdoor threat that has already been implanted into the model. There are two main strategies for post-training defense. (1) Pruning-based Methods [33, 56]: These approaches focus on identifying and removing the most suspicious neurons. (2) Fine-tuning-based Methods [53, 68, 39, 29]: These methods purify the backdoored model by directly tuning its parameters with clean data.
@@
Backdoor attacks and defenses on CLIP. Although vision-language models (VLMs) [45, 23, 32] have demonstrated significant improvements across various fields, recent research [4, 62] has revealed vulnerabilities of contrastively pre-trained VLMs to backdoor attacks, drawing substantial attention in the field of backdoor attacks and defenses on CLIP-like models. In terms of attacks, BadCLIP [31] optimizes visual trigger patterns within a dual-embedding guided framework, rendering the attack effective and undetectable. Moreover, many backdoor attack methods on multimodal large language models [35, 36, 30, 42, 37] have been proposed recently. Defense strategies can be primarily categorized into two categories: pre-training defense and post-training defense. For pre-training defense [61, 22, 60], SAFECLIP [61] prevents learning from poisoned examples by dynamically dividing the training data into the safe and risky sets by cosine similarity and only applying uni-modal contrastive loss to different modalities of the risky set. For fine-tuning defense [3, 59, 25], CleanCLIP [3] reduces the impact of spurious relationships of backdoor attacks by enforcing the model to realign representations for independent modalities.
@@
Prompt learning for vision-language models.  Lately, various fine-tuning methods have been introduced to further enhance the impressive capabilities of vision-language models (VLMs) across diverse downstream tasks [26, 20, 67, 58], with prompt learning [66, 46, 64] emerging as a particularly efficient and effective approach. Specifically, CoOp [66] first introduces prompt learning to VLMs and proposes a method that optimizes the context of the text prompt, e.g., “a photo of a <CLS>.” while keeping the class name tokens fixed. VPT [24] adjusts the visual features by concatenating the image or image features with the learnable prompt. Besides, BadCLIP [1] injects backdoor during the prompt learning stage, enabling the malicious alteration of text features through trigger-aware prompts for a powerful attack. We notice that despite the effectiveness of prompt learning for enhancing the model’s robustness [25, 27, 8, 65, 11], limited research has explored its use in defending against multimodal backdoor attacks. In particular, current post-training backdoor defense methods for VLMs mainly focus on fine-tuning the whole backdoored models, overlooking the potential of prompt learning. To address the gap, we propose Repulsive Visual Prompt Tuning, which incorporates learnable parameters into features to defend against multimodal backdoor attacks while freezing the model’s parameters.
@@ 3 The Proposed Approach
In this section, we first outline the basic settings of the threat model and defender’s goal in Part 3.1, followed by a detailed illustration of how the adversary launches backdoor attacks on CLIP in Part 3.2. Eventually, in Part 3.3, we present Repulsive Visual Prompt Tuning, designed to defend backdoored CLIP against multiple multimodal backdoor attacks.
@@ 3.1 Basic Settings
Threat model.  We assume the adversary has access to the training dataset and can produce a backdoored model by injecting carefully crafted poisoned data into the original clean dataset. For multimodal attacks, the poisoned data usually consists of images with the trigger pattern and the proxy captions for the target class. Once trained on the poisoned dataset, the model will misclassify inputs containing the trigger pattern into the predefined target class while maintaining correct classification for all other benign inputs.
@@ 3.2 Backdoor Attacks on CLIP
Upon training with $\widetilde{\mathcal{D}}$, the encoders of CLIP will be backdoored, denoted as $\{\widetilde{\mathcal{V}}(\cdot),\widetilde{\mathcal{T}}(\cdot)\}$. For the backdoored version of CLIP, the trigger pattern $\Theta$ will have a spurious correlation with the proxy caption of the target class $y_{t}$. Thus, during the inference stage, when the model encounters the attacked images containing the trigger pattern, the predicted posterior probability of the image being classified as the target class $y_{t}$ will increase significantly. In the meantime, the backdoored CLIP performs normally on benign inputs.
@@ 3.3 Repulsive Visual Prompt Tuning
Formally, assume the defender is given the backdoored CLIP, denoted by $\{\widetilde{\mathcal{V}}(\cdot),\widetilde{\mathcal{T}}(\cdot)\}$ alongside a small set of clean labeled samples $\{\boldsymbol{x}_{i},y_{i}\}^{N}_{i=1}$, $y_{i}\in\mathcal{Y}=\{1,\dots,C\}$. In particular, our focus is on CLIP with the Vision Transformer [12] as its visual encoder. For clean input $\boldsymbol{x}_{i}$, the feature representation at the $d$-th layer of CLIP’s visual encoder is $\boldsymbol{f}^{d}_{i}=[\boldsymbol{c}^{d}_{i},\boldsymbol{e}^{d}_{i}]$, where $\boldsymbol{c}^{d}_{i}\in\mathbb{R}^{1\times d_{v}}$ denotes the class embedding that will finally be projected to the vision-language joint space and $\boldsymbol{e}^{d}_{i}\in\mathbb{R}^{M\times d_{v}}$ are the patch embeddings. Specifically, $M$ and $d_{v}$ are hyperparameters representing the number of patches and dimensions of the visual embedding.
@@
$\mathcal{L}_{\mathrm{CE}}=-\sum_{i=1}^{N_{b}}\nolimits\mathrm{log}(\frac{\widetilde{\phi}(\boldsymbol{x}_{i},T(y_{i}))/\tau}{\sum_{c=1}^{C}\widetilde{\phi}(\boldsymbol{x}_{i},T(c))/\tau}),$ (4)
where $\tau$ denotes a temperature parameter to control the logit distribution and is set as specified in [45], $T(\cdot)$ is the proxy caption of the input class, e.g., “a photo of a <CLS>.” and $\widetilde{\phi}(\boldsymbol{x}_{i},\boldsymbol{t}_{j})=\cos(\boldsymbol{P}_{\mathrm{I}}\widetilde{\mathcal{V}}(\boldsymbol{x}_{i}),\boldsymbol{P}_{\mathrm{T}}\widetilde{\mathcal{T}}(\boldsymbol{t}_{j}))$ means cosine similarity of the outputs from the recovering encoders.
@@
For further explanation, RVPT can be interpreted through the lens of natural selection [9]. To be specific, FR loss creates a challenging environment for visual prompt tuning to optimize CE loss. In order to simultaneously minimize CE loss and maximize the discrepancy between the prompted features and original features, the visual prompts must learn to select the most suitable features to optimize CE loss while ignoring those that are non-essential. Ultimately, this process leads to a more compact visual encoding mechanism, enhancing CLIP’s perturbation resistivity to trigger patterns.

### v3

@@ Abstract
Multimodal contrastive learning models (e.g., CLIP) can learn high-quality representations from large-scale image-text datasets, while they exhibit significant vulnerabilities to backdoor attacks, raising serious safety concerns. In this paper, we reveal that CLIP’s vulnerabilities primarily stem from its tendency to encode features beyond in-dataset predictive patterns, compromising its visual feature resistivity to input perturbations. This makes its encoded features highly susceptible to being reshaped by backdoor triggers. To address this challenge, we propose Repulsive Visual Prompt Tuning (RVPT), a novel defense approach that employs deep visual prompt tuning with a specially designed feature-repelling loss. Specifically, RVPT adversarially repels the encoded features from deeper layers while optimizing the standard cross-entropy loss, ensuring that only predictive features in downstream tasks are encoded, thereby enhancing CLIP’s visual feature resistivity against input perturbations and mitigating its susceptibility to backdoor attacks. Unlike existing multimodal backdoor defense methods that typically require the availability of poisoned data or involve fine-tuning the entire model, RVPT leverages few-shot downstream clean samples and only tunes a small number of parameters. Empirical results demonstrate that RVPT tunes only 0.27% of the parameters in CLIP, yet it significantly outperforms state-of-the-art defense methods, reducing the attack success rate from 89.70% to 2.76% against the most advanced multimodal attacks on ImageNet and effectively generalizes its defensive capabilities across multiple datasets.
@@ 1 Introduction
Figure 1: Perturbation Resistivity (PR) across different layers of the encoders under various perturbations, including the trigger pattern. A higher PR value indicates less sensitivity to input perturbation. Four encoder variants are evaluated on ImageNet: (a) ViT in backdoored CLIP radford2021clip . (b) ViT in backdoored CLIP tuned with Visual Prompt Tuning (VPT) on ImageNet. (c) ViT of the backdoored CLIP tuned with our RVPT on ImageNet. (d) ViT trained exclusively on clean ImageNet paszke2019pytorch . Specifically, CLIP is backdoored by the trigger of BadCLIP liang2024badclip . Detailed experimental settings and additional PR results for CLIP backdoored by other triggers can be found in Appendix A.
@@
Recently, Contrastive Language-Image Pretraining (CLIP) radford2021clip has emerged as a powerful base model across multiple domains. Unlike traditional models that rely on uni-modal supervision, CLIP aligns visual and textual signals to learn rich representations from numerous image-text pairs in the pre-training stage. This unique training paradigm enables its impressive performance in open-vocabulary visual recognition wu2024towards and robustness to distribution shift fang2022data .
@@
However, despite CLIP’s success in visual recognition, recent research carlini2021poisoning ; 10646610 has revealed its vulnerabilities against backdoor attacks in downstream tasks. The vulnerabilities can be roughly categorized into two aspects. First, concerns arise from its training data collection process: the data is noisy, uncurated, and sourced from the web, which facilitates adversarial poisoning of the image-text pairs. Second, CLIP demonstrates heightened susceptibility to a smaller number of poisoned training samples: recent studies carlini2021poisoning have empirically shown that poisoning CLIP may require injecting orders of magnitude fewer poisoned data into the training dataset than poisoning conventional supervised models. Once poisoned during pre-training, the backdoored CLIP will inevitably classify images with an adversarially designed trigger into the predefined target class.
@@
CLIP’s vulnerabilities to backdoor attacks have inspired the development of various defense methods. However, most of these methods necessitate fine-tuning the entire model’s parameters bansal2023cleanclip or require the availability of poisoned data yang2023safeclip ; ishmam2024semantic ; yang2023robust , which could be resource-intensive or even unrealistic. Thus, to achieve an efficient, effective, and practical defense, we are dedicated to developing a defense method based on visual prompt tuning (VPT) jia2022visual , utilizing only a small portion of downstream clean samples to encourage the model to ignore the trigger feature in the poisoned images. However, simply adopting VPT proves ineffective experimentally (see Figure 1(b)), as the trigger feature cannot be extracted from clean samples, preventing VPT from learning how to suppress them.
@@
In this paper, we find that CLIP’s visual embeddings remain highly sensitive to small input perturbations even after VPT, leaving it susceptible to potential backdoor attacks. To quantify this sensitivity, we introduce the Perturbation Resistivity (PR) metric, defined as the cosine similarity between the visual embedding $f^{v}(\boldsymbol{x})$ of an image $\boldsymbol{x}$ and that of its perturbed counterpart: $\boldsymbol{x}+\boldsymbol{\delta}$:
$\text{PR}(\boldsymbol{x},\boldsymbol{\delta})=\cos(f^{v}(\boldsymbol{x}),f^{v}(\boldsymbol{x}+\boldsymbol{\delta})),$ (1)
where $f^{v}(\cdot)$ denotes the visual feature extractor. A lower PR value implies larger visual embedding drift due to input perturbations, hence greater sensitivity and higher susceptibility to backdoor triggers. To illustrate this vulnerability of CLIP, we apply a suite of perturbations to ImageNet samples and plot the layer-wise mean PR for several model variants. The results in Figure 1 show that the deeper layers of standard CLIP (Figure 1(a)) and standard CLIP tuned with VPT (Figure 1(b)) exhibit significantly lower PR than a ViT trained exclusively on clean ImageNet (Figure 1(d)), confirming that CLIP’s embeddings are easily distorted. Because a backdoor trigger acts as a structured perturbation, this low PR allows the trigger to maliciously reshape the encoded visual features layer by layer, enabling a successful attack. We therefore regard PR as a proxy for backdoor robustness, which means the backdoor-robust CLIP should have a higher PR than backdoor-vulnerable CLIP (i.e., the standard CLIP). Our analysis suggests that CLIP’s low PR originates from its large-scale, noisy pre-training dataset: the encoder captures many off-dataset, non-predictive features, which magnify the impact of perturbations. In contrast, the ImageNet-supervised ViT only learns in-dataset predictive features of ImageNet, prohibiting it from encoding irrelevant features beyond ImageNet and achieving a much higher PR. Hence, an effective defense (ours shown in Figure 1(c)) should suppress the irrelevant features CLIP inherently captures and guide it to focus on in-dataset predictive patterns, thereby enhancing both its perturbation resistivity and backdoor robustness.
@@
Inspired by this finding, we propose a novel approach termed Repulsive Visual Prompt Tuning (RVPT), which improves CLIP’s PR by making it exclusively encode the predictive features in the downstream tasks in a few-shot manner. As illustrated in Figure 2, RVPT jointly optimizes two objectives: (1) a multi-layer feature-repelling loss that minimizes the cosine similarity between prompted features and original CLIP features at each layer, and (2) cross-entropy loss for downstream task accuracy. This mechanism forces the model to only extract features that contribute to CE loss optimization during training, while ignoring non-essential features (e.g., noise or trigger pattern). As a result, Figure 1(c) shows that RVPT significantly improves the PR of CLIP on ImageNet, hence effectively defending against the state-of-the-art multimodal backdoor attack. Moreover, extensive experiments demonstrate our method’s effectiveness across multiple datasets and backdoor attacks.
@@ 2 Related Work
Backdoor attacks and defenses on supervised learning. Backdoor attacks have become an increasingly serious security issue because more and more practitioners choose to adopt third-party datasets, platforms, or even backbones to reduce costs. Research on backdoor attacks has primarily focused on designing triggers that enhance stealthiness chen2017blend ; turner2019labelconsistent and effectiveness optba ; nguyen2020input . For backdoor defenses, researchers employ various strategies to mitigate backdoor attacks, including: (1) pre-processing defense tran2018spectral (2) pre-training defense chen2022effective (3) post-training defense zhu2024neural ; MM-BD (4) test-time defense feng2023detecting . Specifically, our work aligns with the post-training defense category, which assumes that defenders do not have access to the poisoned training data and must eliminate the backdoor threat that has already been implanted into the model. There are two main strategies for post-training defense. (1) Pruning-based Methods liu2018fine-pruning ; wu2021adversarial : These approaches focus on identifying and removing the most suspicious neurons. (2) Fine-tuning-based Methods neuralcleanse ; zhu2024neural ; min2023towards ; li2021neural : These methods purify the backdoored model by directly tuning its parameters with clean data.
@@
Backdoor attacks and defenses on vision-language models. Although vision-language models (VLMs) radford2021clip ; pmlr-v139-align ; liu2023llava have demonstrated significant improvements across various fields, recent research carlini2021poisoning ; yang2023data has revealed vulnerabilities of contrastively pre-trained VLMs to backdoor attacks, drawing substantial attention in the field of backdoor attacks and defenses on CLIP-like models. In terms of attacks, BadCLIP liang2024badclip optimizes visual trigger patterns within a dual-embedding guided framework, rendering the attack effective and undetectable. Moreover, many backdoor attack methods on multimodal large language models lu2024test ; lyu2024trojvlm ; liang2024revisiting ; ni2024physical ; lyu2024backdooring ; liang2025vl ; yuan2025badtoken ; liu2025stealthy have been proposed recently. Defense strategies can be primarily categorized into two categories: pre-training defense and post-training defense. For pre-training defense yang2023safeclip ; ishmam2024semantic ; yang2023robust , SAFECLIP yang2023safeclip prevents learning from poisoned examples by dynamically dividing the training data into the safe and risky sets by cosine similarity and only applying uni-modal contrastive loss to different modalities of the risky set. For fine-tuning defense bansal2023cleanclip ; xun2024ta ; kuang2024adversarial , CleanCLIP bansal2023cleanclip reduces the impact of spurious relationships of backdoor attacks by enforcing the model to realign representations for independent modalities. There are also some work focusing on test-time defense on multimodal backdoor attacks niu2024bdetclip ; hecloser ; huang2025detecting .
@@
Prompt learning for vision-language models. Lately, various fine-tuning methods have been introduced to further enhance the impressive capabilities of vision-language models (VLMs) across diverse downstream tasks LIXWH ; huang2024froster ; zhou2023anomalyclip ; wu2024clipselfvisiontransformerdistills , with prompt learning zhou2022coop ; rao2022denseclip ; zhang2024candidate emerging as a particularly efficient and effective approach. Specifically, CoOp zhou2022coop first introduces prompt learning to VLMs and proposes a method that optimizes the context of the text prompt, e.g., “a photo of a <CLS>.” while keeping the class name tokens fixed. VPT jia2022visual adjusts the visual features by concatenating the image or image features with the learnable prompt. Besides, BadCLIP bai2023badclip injects backdoor during the prompt learning stage, enabling the malicious alteration of text features through trigger-aware prompts for a powerful attack. We notice that despite the effectiveness of prompt learning for enhancing the model’s robustness kuang2024adversarial ; li2024one ; chuang2023debiasingvisionlanguagemodelsbiased ; zheng2022prompt ; dong2023lptlongtailedprompttuning , limited research has explored its use in defending against multimodal backdoor attacks. In particular, current post-training backdoor defense methods for VLMs mainly focus on fine-tuning the whole backdoored models, overlooking the potential of prompt learning. To address the gap, we propose RVPT, which incorporates learnable parameters into features to defend against multimodal backdoor attacks while freezing the model’s parameters.
@@ 3 The Proposed Approach
Overview. In this section, we first outline the basic settings of the threat model and defender’s goal in Section 3.1, followed by a detailed illustration of how the adversary launches backdoor attacks on CLIP in Section 3.2. Eventually, in Section 3.3, we present Repulsive Visual Prompt Tuning, designed to defend backdoored CLIP against multiple multimodal backdoor attacks.
@@ 3.1 Basic Settings
Threat model. We assume the adversary has access to the training dataset and can produce a backdoored model by injecting crafted poisoned data into the original dataset. For multimodal attacks, the poisoned data usually consists of images with the trigger pattern and the proxy captions for the target class. Once trained on the poisoned dataset, the model will misclassify inputs with the trigger pattern into the predefined target class while maintaining classify correctly for the benign inputs.
@@ 3.2 Backdoor Attacks on CLIP
Upon training with $\widetilde{\mathcal{D}}$, the encoders of CLIP will be backdoored, denoted as $\{\widetilde{\mathcal{V}}(\cdot),\widetilde{\mathcal{T}}(\cdot)\}$. For the backdoored version of CLIP, the trigger pattern $\Theta$ will have a spurious correlation with the proxy caption of the target class $y_{t}$. Thus, during the inference stage, when the model encounters the images containing the trigger pattern, its prediction on the target class $y_{t}$ will increase significantly. In the meantime, the backdoored CLIP performs normally on benign inputs.
@@ 3.3 Repulsive Visual Prompt Tuning
Formally, assume the defender is given the backdoored CLIP, denoted by $\{\widetilde{\mathcal{V}}(\cdot),\widetilde{\mathcal{T}}(\cdot)\}$ alongside a small set of clean labeled samples $\{\boldsymbol{x}_{i},y_{i}\}^{N}_{i=1}$, $y_{i}\in\mathcal{Y}=\{1,\dots,C\}$. In particular, our focus is on CLIP with the Vision Transformer dosovitskiy2020vit as its visual encoder. For clean input $\boldsymbol{x}_{i}$, the feature representation at the $d$-th layer of CLIP’s visual encoder is $\boldsymbol{f}^{d}_{i}=[\boldsymbol{c}^{d}_{i},\boldsymbol{e}^{d}_{i}]$, where $\boldsymbol{c}^{d}_{i}\in\mathbb{R}^{1\times d_{v}}$ denotes the class embedding that will finally be projected to the vision-language joint space and $\boldsymbol{e}^{d}_{i}\in\mathbb{R}^{M\times d_{v}}$ are the patch embeddings. Specifically, $M$ and $d_{v}$ are hyperparameters representing the number of patches and dimensions of the visual embedding.
@@
$\mathcal{L}_{\mathrm{CE}}=-\sum_{i=1}^{N_{b}}\nolimits\mathrm{log}(\frac{\exp(\widetilde{\phi}(\boldsymbol{x}_{i},T(y_{i}))/\tau)}{\sum_{c=1}^{C}\exp(\widetilde{\phi}(\boldsymbol{x}_{i},T(c))/\tau)}),$ (4)
where $\tau$ denotes a temperature parameter to control the logit distribution and is set as specified in radford2021clip , $T(\cdot)$ is the proxy caption of the input class, e.g., “a photo of a <CLS>.” and $\widetilde{\phi}(\boldsymbol{x}_{i},\boldsymbol{t}_{j})=\cos(\boldsymbol{P}_{\mathrm{I}}\widetilde{\mathcal{V}}(\boldsymbol{x}_{i}),\boldsymbol{P}_{\mathrm{T}}\widetilde{\mathcal{T}}(\boldsymbol{t}_{j}))$ means cosine similarity of the outputs from the recovering encoders.
@@
For further explanation, RVPT can be interpreted through the lens of natural selection article . To be specific, FR loss creates a challenging environment for visual prompt tuning to optimize CE loss. In order to simultaneously minimize CE loss and maximize the discrepancy between the prompted features and original features, the visual prompts must learn to select the most suitable features to optimize CE loss while ignoring those that are non-essential. Ultimately, this process leads to a more compact visual encoding mechanism, enhancing CLIP’s perturbation resistivity to trigger patterns.